OpenStudy (anonymous):

to solve by completing the square what value should you add to each side of the equation x^2+20x=-8

OpenStudy (campbell_st):

a perfect square is \[(x +b)^2 = x^2 + 2bx + b^2\] so halve to middle term, then square it... then add the value to both sides of the equation.

OpenStudy (wolf1728):

To complete the square we have to have the x² term have no exponent and move the non-x term to the right side of the equation.(The equation was already given this way). Then we take the coefficient of x (20) which we divide by 2 square then add it to both sides 20/2 = 10 and 10² = 100 x^2 + 20x + 100 = 92 (x+10) * (x+10) = 92 taking the square root of both sides x +10 = sqrt(92) x = 9.5916630466 -10 x = -0.4083369534

OpenStudy (campbell_st):

just a quick point... you will have 2 solutions to your problem not 1 as you are looking at \[x + 10 = \pm \sqrt{92}\]

OpenStudy (wolf1728):

campbell_st Unbelievable - I went to the trouble of typing a rather detailed explanation of completing the square - and I only gave half the answer. Good catch!!